# Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology

### v2

@@ 3.2 Proposed methods
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$, and the preprocessed matrix is
@@
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements in the original coefficient matrix $A$ closer to the main diagonal elements. Even if the matrix structure is closer to block diagonal form and $P$ is a permutation matrix (i.e., with $P^{T}P=E$), the linear system (1.1) becomes
@@
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate block diagonal matrix, as long as $\widetilde{A}$ is appropriately partitioned, orthogonality between blocks can be achieved, see Figure 3.
Figure 3:   The orthogonality representation of the matrix Blackhole after being rearranged using the RCM algorithm.
Consider dividing $\widetilde{A}$ into $k$ blocks, taking $\left[m/k\right]$ elements from each block, and storing the remaining elements in the $k+1$-th block. Perform corresponding partitioning on $\widetilde{f}$, that is, partitioning:
$\widetilde{A}=\begin{bmatrix}A_{\tau_{1}}\\ A_{\tau_{2}}\\ ...\\ A_{\tau_{k}}\\ A_{\tau_{k+1}}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{\tau_{1}}\\ f_{\tau_{2}}\\ ...\\ f_{\tau_{k}}\\ f_{\tau_{k+1}}\\ \end{bmatrix}.$
@@
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. So obviously, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non orthogonal hyperplanes. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
@@
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{\tau_{i}}$ as $\overline{A}_{\tau_{i}}$, $i=1,2,...,k+1$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates to form an inner product table $G$, $G\left(i,j\right)$ which is used to measure the orthogonality between blocks $\overline{A}_{\tau_{i}}$ and $\overline{A}_{\tau_{j}}$. Moreover, as $\widetilde{A}$ is an approximate block diagonal matrix, it is unreasonable and convenient to use the inner product of centroid coordinates to reflect the orthogonality between blocks. And we divide the matrix $\widetilde{A}$ into $k+1$ blocks in advance, in fact, we only need to do inner product on the centroid coordinates of these $k+1$ blocks. Moreover, since $G$ is a symmetric matrix, we only need to calculate $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times, because $k\ll m$, this is a relatively small number.
Next, based on our above analysis, our aim is to group two row blocks that are mutually orthogonal, i.e., $G\left(i,j\right)=0$ into one class to accelerate the Kaczmarz method. However, it is evident that there may be more than one block orthogonal to any block, which can easily result in some row blocks being in multiple categories simultaneously. In order to avoid this situation, we stipulate that for each block $A_{\tau_{i}}$, $i=1,2,..k$, we select the first block in $\left\{A_{\tau_{i}},A_{\tau_{i+1}},...,A_{\tau_{k+1}}\right\}$ that has an inner product of 00 with the centroid coordinates of the current block $A_{\tau_{i}}$ and place it in one category, marking the current class as $class\left\{{i}\right\}$, and ensuring that the blocks between each class are not duplicated. For blocks that have not entered any class, in order to avoid non-convergence caused by information loss, we classify all these blocks into a new class, denoted as $class\left\{l\right\}$. This regulation is reasonable because $\widetilde{A}$ approximates the properties of block diagonal matrices. Obtain solutions through block Kaczmarz iteration method on two categories
@@
1:sparse matrix $A$, right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, the maximum number of iterations $l$.
@@
4:Apply the k-means method to $\widetilde{A}$ to obtain a partition $\{\tau_{1},...,\tau_{k+1}\}$ of the row $\{1,...m\}$, and obtain $k+1$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k+1$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
5:Calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates of each block to form the inner product table $G$.
6:According to the above regulations, place mutually orthogonal blocks in $class\left\{i\right\}$, and ensure that the blocks between each class are not duplicated. Divide all remaining blocks into a new $class\left\{l\right\}$.
@@
8:     Select   $\tau_{j1},\tau_{j2}\in class\left\{{i}\right\}$
@@
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right)$
$\widetilde{y}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{j2}}^{\dagger}\left(f_{\tau_{j2}}-A_{\tau_{j2}}\widetilde{x}_{{j+1}/2}\right)$
10:     Select   $\tau_{j}\in class\left\{{l}\right\}$
@@
$\widetilde{x}_{j+1}=\widetilde{y}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{y}_{j+1}\right)$
@@
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times inner product and then orthogonal partitioning largely avoids this problem. On the one hand, the computation time of inner product is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
@@ 4 Convergence analysis
Lemma 1.19 Assumes that a system of linear equations with a coefficient matrix of $\widetilde{A}\in{{\mathbb{R}}^{m\times m}}$ and a right-hand vector of $\widetilde{f}\in{{\mathbb{R}}^{m}}$ is compatible. Using the idea of uniform partitioning, the row vectors of the coefficient matrix $\widetilde{A}$ are partitioned to obtain $A_{\tau_{1}},A_{\tau_{2}},...,A_{\tau_{k+1}}$. Singular value decomposition is performed on the submatrix $A_{\tau_{i}}\in{{\mathbb{R}}^{m_{\tau_{i}}\times m}}$
$A_{\tau_{i}}=U_{\tau_{i}}\Sigma_{\tau_{i}}V_{\tau_{i}}^{H}$ (4.2)
@@
where $U_{\tau_{i}}$ is a $m_{\tau_{i}}\times m_{\tau_{i}}$ unitary matrix. $\Sigma_{\tau_{i}}$ is a $m_{\tau_{i}}\times m$ diagonal matrix. $V_{\tau_{i}}^{H}$, the conjugate transpose of $V_{\tau_{i}}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{\tau_{i}}$ are singular values of $A_{\tau_{i}}$. Then $A_{\tau_{i}}^{\dagger}=V_{\tau_{i}}\Sigma_{\tau_{i}}^{H}U_{\tau_{i}}^{H}$, and $A_{\tau_{i}}^{\dagger}A_{\tau_{i}}=V_{\tau_{i}}MV_{\tau_{i}}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{\tau_{i}}$.
Lemma 2. Assumes that $V$ is a unitary matrix of $m\times m$ order, which can be considered as a combination of column vectors, denoted as $V=[v_{1},v_{2},...,v_{m}]$ then
(1) $v_{i}^{H}v_{j}=\left\{\begin{aligned} 0,~{}~{}~{}~{}~{}i\neq j\\ 1,~{}~{}~{}~{}~{}i=j\end{aligned}\right.$
(2) For any $m$ dimensional vector $u$, it can be uniquely represented as a column linear combination $u=\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}$, and the Euclidean norm $\left\|u\right\|_{2}^{2}=\left\|\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}\right\|_{2}^{2}=\sum_{i=1}^{m}\alpha_{i}^{2}$.
Similar to Ref. 20 , we have the following convergence theorem.
@@
Theorem 1.For the linear system (1.1) $Ax=f$, the iteration sequence $\left\{x_{j}\right\}_{j=0}^{\infty}$ generated by POBK algorithm GREBK(k) converges to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
Proof. By POBK algorithm GREBK(k), the main calculation is for block wise linear systems in the following form
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right).$
For convenience, denote $A_{\tau_{j1}}=A_{\tau_{0}}$. For the submatrix $A_{\tau_{0}}$ of $\widetilde{A}$, according to Lemma 1, there are $A_{\tau_{0}}=U_{\tau_{0}}\Sigma_{\tau_{i}}V_{\tau_{0}}^{H}$, $A_{\tau_{0}}^{\dagger}=V_{\tau_{0}}\Sigma_{\tau_{i}}^{H}U_{\tau_{0}}^{H}$, and therefore $A_{\tau_{0}}^{\dagger}A_{\tau_{0}}=V_{\tau_{0}}MV_{\tau_{0}}^{H}$, $M=diag\left(1,...,1,0,...,0\right)$, where the number of 1 is the rank of $A_{\tau_{0}}$, then
$x_{1}-x_{*}$ $=x_{0}+A_{\tau_{0}}^{\dagger}\left(f_{\tau_{0}}-A_{\tau_{0}}x_{0}\right)-x_{*}$ (4.3)
$=x_{0}-x_{*}+A_{\tau_{0}}^{\dagger}\left(A_{\tau_{0}}x_{*}-A_{\tau_{0}}x_{0}\right)$
$=\left(I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\right)\left(x_{0}-x_{*}\right).$
Due to $I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}$ being an orthogonal projection operator and $x_{0}-x_{*}$ being linearly represented by the column vectors of $V_{\tau_{0}}$, i.e.,
$x_{0}-x_{*}=\alpha_{\tau_{0}1}v_{\tau_{0}1}+\alpha_{\tau_{0}2}v_{\tau_{0}2}+...+\alpha_{\tau_{0}m}v_{\tau_{0}m}=V_{\tau_{0}}\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}.$ (4.4)
So
$\left\|x_{1}-x_{*}\right\|_{2}^{2}$ $=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$ (4.5)
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|V_{\tau_{0}}MV_{\tau_{0}}^{H}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|M\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|\left[\alpha_{\tau_{0}1},...,\alpha_{\tau_{0}r},0,...,0\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\gamma_{\tau_{0}}\left\|x_{0}-x_{*}\right\|_{2}^{2}$
$=\left(1-\gamma_{\tau_{0}}\right)\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}$
where $\gamma_{\tau_{0}}=\sum_{i=1}^{r}\left(\alpha_{\tau_{0}i}\right)^{2}/\sum_{i=1}^{m}\left(\alpha_{\tau_{0}i}\right)^{2}$, obviously $10<\gamma_{\tau_{0}}<$ 1. Therefore
$\centering\|x_{j}-x_{\star}\|_{2}^{2}=(1-\gamma_{j})\|x_{j-1}-x_{\star}\|_{2}^{2}<\|x_{j-1}-x_{\star}\|_{2}^{2}.\@add@centering$ (4.6)
i.e., $\mathop{\lim}\limits_{j\to\infty}{\left\|{{x_{j}}-{x_{\star}}}\right\|_{2}}=0$. POBK algorithm GREBK(k) converges. ∎
@@
For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes see Figure 4:
@@
Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above situations, we have that
@@
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left(1-\gamma_{\tau_{0}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.9)
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the hyperplane angle $\theta$. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between the two blocks in $class\left\{i\right\}$ is close to 90 degrees, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $class\left\{l\right\}$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate diagonal matrix, the blocks in the matrix have better orthogonality than the original matrix $A$, and the convergence rate is also faster than the GRBK, RBK (k), and GREBK (k) algorithms.

### v3

@@ 3.2 The Proposed methods
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$ as follows.
@@
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements of the original coefficient matrix $A$ closer to the main diagonal, i.e., it makes the structure of matrix $A$ closer to the banded matrix form, and $P$ is a substitution matrix (i.e., there is $P^{T}P=E$), so the linear system (1.1) becomes
@@
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row chunking of $\widetilde{A}$, see Figure 2.
Due to the natural orthogonality of banded matrices (see Figure 2), we directly chunk $\widetilde{A}$ uniformly in order. Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding chunking is done for $f$,
$\widetilde{A}=\begin{bmatrix}A_{1}\\ A_{2}\\ ...\\ A_{k}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{1}\\ f_{2}\\ ...\\ f_{k}\\ \end{bmatrix}.$
@@
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. Then, according to the analysis in Section 3.1,, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non-orthogonal hyperplanes, see Figure 1. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
@@
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $C\left(i,j\right)=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $C\left(i,j\right)$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$. And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks. And we divide the matrix $\widetilde{A}$ into $k$ blocks beforehand, and we only need to calculate the cosine of the angle between these $k$ blocks, and since $C$ is a symmetric matrix, we only need to calculate $\frac{k\left(k-1\right)}{2}$ times in practice. Since $k\ll m$, this is a fairly small number.
Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $C(i,j)=0$, into one class. Here we set a threshold $thr$ such that when $C(i,j)<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal, i.e., so that $C(i,j)=0$. However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time. In order to avoid this situation, we stipulate that for each block $A_{i}$, $i=1,2,..k$, the first block $A_{j}\in\left\{A_{i},A_{i+1},...,A_{k}\right\}$ orthogonal to $A_{i}$ is selected and placed in a class, noting the current class as $class\left\{i\right\}$, and requiring that there is no duplication of blocks between classes. We merge these orthogonal classes, denoted $Oclass$. For the blocks that do not go into $Oclass$, in order to avoid the non-convergence situation caused by the loss of information, we assign all of these blocks in a new class, denoted $Nclass$. This regulation is reasonable because of the nature of the A-approximate banded matrix. The solution is obtained by the block Kaczmarz iterative method on both categories:
@@
1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
@@
4:Perform uniform chunking of $\widetilde{A}$ and $\widetilde{f}$ in sequential order, and obtain $k$ groups $A_{i}$ and $f_{i}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{1},...,\overline{A}_{k}]$, $\overline{f}=[\overline{f}_{1},...,\overline{f}_{k}]$, where $\overline{A}_{i}$ and $\overline{f}_{i}$ respectively represent the center of $A_{i}$ and $f_{i}$.
5:Compute the cosine of the angle $C\left(i,j\right)$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
6:According to the above regulations, place all two-by-two mutually orthogonal blocks in a category $Oclass$, place all remaining blocks in a new class $Nclass$.
@@
8:     Select   $\tau_{1},\tau_{2}\in Oclass$
@@
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{1}}^{\dagger}\left(f_{\tau_{1}}-A_{\tau_{1}}\widetilde{x}_{j}\right)$
$\widetilde{x}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{2}}^{\dagger}\left(f_{\tau_{2}}-A_{\tau 2}\widetilde{x}_{{j+1}/2}\right)$
10:     Select   $\tau_{j}\in Nclass$
@@
$\widetilde{x}_{j+2}=\widetilde{x}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j+1}\right)$
@@
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{k\left(k-1\right)}{2}$ calculations of the angle cosine values and then orthogonal chunking avoids this problem to a large extent.On the one hand, the computation time of the angle cosine value is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
@@ 4 Convergence analysis
Lemma 1 [25]. Assumes that the singular value decomposition is performed on the submatrix $A_{i}\in{{\mathbb{R}}^{m_{i}\times m}}$,
$A_{i}=U_{i}\Sigma_{i}V_{i}^{H}$ (4.2)
@@
where $U_{i}$ is a $m_{i}\times m_{i}$ unitary matrix. $\Sigma_{i}$ is a $m_{i}\times m$ diagonal matrix. $V_{i}^{H}$, the conjugate transpose of $V_{i}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{i}$ are singular values of $A_{i}$. Then the Moore–Penrose pseudoinverse of $A_{i}$ $A_{i}^{\dagger}=V_{i}\Sigma_{i}^{H}U_{i}^{H}$, and $A_{i}^{\dagger}A_{i}=V_{i}MV_{i}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{i}$.
Lemma 2 [13]. For any $A\in{{\mathbb{R}}^{n\times m}}$, we have
$\sigma_{min}(A^{\dagger})=\frac{1}{\sigma_{max}(A)},$
where $\sigma_{min}\left(.\right)$ denotes the smallest nonzero singular value of a matrix and $\sigma_{max}\left(.\right)$ denotes the largest singular value of a matrix.
Similar to Ref. [16], we have the following convergence theorem.
@@
Theorem 1. If the linear system (1.1) $Ax=f$ is consistent, then the POBK algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$. In particular, we have the following linear convergence rate in expectation:
$\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{jk}\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.3)
Proof. When $\tau_{j}\in Nclass$, the POBK algorithm involves one block per iteration and computes the linear system by the following form
${x}_{j+1}={x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}{x}_{j}\right).$ (4.4)
According to Lemma 1, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$. Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
$x_{j+1}-x_{*}$ $=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}$ (4.5)
$=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)$
$=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).$
Therefore, we have
$\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$ (4.6)
$=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.$
Taking the conditional expectation on both sides of inequality (4.6), since the matrix $\widetilde{A}$ is uniformly partitioned, the probability of each block being selected is equal, i.e., $p_{\tau}=\frac{1}{k}$, we get
$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]$ $=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$ (4.7)
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$
$\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}$
$\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},$
and assuming that $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.8)
By taking the full expectation on both sides of the above equation, we have
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.9)
When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the POBK algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}$ (4.10)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
We divide the matrix into $k$ blocks, where $k_{1}$ blocks are partitioned into $Oclass$ and $k_{2}$ blocks are partitioned into $Nclass$, and $k_{1}+k_{2}=k$. According to our POBK algorithm, iterating first in $Oclass$ and then in $Nclass$, then it is clear that the iteration is done once and
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}$ (4.11)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
The inequality (4.3) follows by induction on $j$. ∎
Remark 1. As can be seen from the above proof, the POBK algorithm converges as expected and the convergence rate is related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks. The power of the convergence rate of the POBK algorithm is much larger than the power of the general Kaczmarz algorithm, and theoretically there will be a much faster convergence speed. In order to visualize the convergence speed of the POBK algorithm more intuitively, below we explain the reason why the POBK algorithm converges faster from the perspective of the hyperplane pinch angle.
@@
For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes, see Figure 4.
@@
Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above cases, we have that
@@
$\left\|x_{j}-x_{*}\right\|_{2}^{2}\leq\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.14)
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the value of the cosine of the hyperplane pinch angle. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between two blocks that are mutually orthogonal in $Oclass$ is close to $90^{\circ}$, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $Nclass$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate banded matrix, the orthogonality between the blocks in the matrix $\widetilde{A}$ is better than the original matrix $A$. Meanwhile, two nearly orthogonal blocks in $Oclass$ have a high probability of being non-adjacent, then the matrices in $Nclass$, although not orthogonal to each other, will have a much lower probability of being adjacent to each other, i.e., the number of times to compute the cosine value of the smaller pinch angle will be much lower, and according to Equation (4.14), its convergence speed is also faster than that of the GRBK, RBK (k), GREBK (k) and aRBK algorithms.